# 2]Centro de Matemática e Aplicações da Universidade da Beira Interior, Rua Marquês D’Ávila e Bolama 6200-001 Covilhã, Portugal

### v1

@@ 4 Characterization of self-resonance
Figure 1: Curvature perturbations for (a) T-model and (b) E-model evaluated 5 e-folds after the end of inflation. Numerical computation is shown in continuous black and Floquet theory as described in this work in dashed green. The vertical grey line marks the scale that exits the horizon at the end of inflation, $k_{\text{end}}$. For comparison, black dotted curves show the numerical computation of curvature perturbations at the end of inflation.
Here we present the Floquet’s exponents for both T- and E-models and compare them with numerical results. These findings can be generalized to any potential whose expansion around $\phi=0$ matches either the expansion of a T- or an E-model.
@@ 4.1 T-model
In this case we have that $\lambda_{3}=0$ and thus $q_{1}=p_{1}=0$, which greatly simplifies the expression for the Floquet exponent
@@
For the rest of this work, we set $\alpha=10^{-2.5}$ for both T- and E-models. To check the validity of this approximation, we have computed the curvature perturbations numerically, using (11) with Bunch-Davies initial conditions and (19). Then, we have used the Floquet theory derived here. Both methods are evaluated 5 e-folds after the end of inflation. Results are displayed in Fig. 1(a) and it shows a good agreement between both methods. The peaks to the right of the highest peak in Fig. 1 correspond to the amplification in higher resonance bands. One could study these bands by considering the frequencies $n=\pm 2,\pm 3\dots$ and their corresponding couplings. However, to do so, one has to increase each time the order in $\epsilon$, increasing the number of harmonics that come into play. We find it sufficient for our purposes to focus on the first resonance band, as it is where the amplification is most pronounced. Finally, in Fig. 2(a) we display $\Re(\mu_{k}^{(T)})$ as a function of $q_{2}$ and $A_{k}$, as well as the trajectories of some modes in the $(q_{2},A_{k})$ plane. We remark that in Fig. 2(a), $p_{2}=0$ and thus the $\mu_{k}^{(T)}$ shown does not correspond to the real time-evolution. However, it is enough to understand why the peak is produced for scales around $k\simeq 10\,k_{\text{end}}$. Those scales cross the $(q_{2},A_{k})$ plane throughout the maximum of $\mu_{k}^{(T)}$. For $k<k_{\text{end}}$, all scales show the same evolution in the $(q_{2},A_{k})$ plane. This is the reason why all of them are amplified in the same amount.
Figure 2: (a) Floquet chart for a T-model with $p_{2}=0$ (just for graphical purposes). White dashed lines correspond to the evolution of the modes in the $(q_{2},A_{k})$ plane and are, from bottom to top: 1, 10, 15, 20, 30, in units of $k_{\text{end}}$. (b) Evolution of $\Re(\mu_{k,E})$ for an E-model as a function of the amplitude of the field and for different values of $k$.
@@ 4.2 E-model
Now we have that $\lambda_{3}\neq 0$ and thus the expression for the Floquet exponent is the full one given in (18), we will call it $\mu_{k}^{(E)}$. Fig. 1(b) displays the curvature perturbations computed using the same method as in Fig. 1(a) (see previous section), this time for the E-model. Despite a numerical factor of $\mathcal{O}(1)$ of discrepancy between the numerics and Floquet theory, we find it enough to explain at which scales occur the maximum amplification. The discrepancy could be because we are neglecting some higher harmonics, which in the case of the E-model could contribute more significantly than in the T-model, as well as more couplings between them. A comparison between Figs. 1(a) and 1(b) reveals a higher amplification for the E-model compared to the T-model, for the same value of $\alpha$. Additionally, the position of the peak of maximum amplification is at higher $k$ for the E-model. Now, regarding the time evolution of $\mu_{k}^{(E)}$ for each $k$-mode, a plot like Fig. 2(a) would not be sufficiently illustrative, given that $\mu_{k}^{(E)}$ depends on 5 parameters. This makes it challenging to understand why the peak is around $k\sim 11\,k_{\text{end}}$. Instead, Fig. 2(b) shows $\Re(\mu_{k}^{(E)})$ as a function of the amplitude of the field for different values of $k$, expressed in units of $k_{\text{end}}$. Here, we observe that for scales around $k\simeq 11\,k_{\text{end}}$, the Floquet exponent exceeds that of other $k$-modes, resulting in higher amplification.
@@ 5 Contrast with quadratic approximation
During preheating, the potential is usually approximated by a parabola ($\lambda_{3}=\lambda=0$), where the solution for the field is just given by $\phi\simeq\phi_{\text{end}}\left(\frac{a_{\text{end}}}{a}\right)^{3/2}\cos(Mt)$ [9]. Considering this and keeping just the dominant terms, (11) is easily reformulated into the following Mathieu equation
$\frac{\mathrm{d}^{2}\tilde{v}_{\bm{k}}}{\mathrm{d}z^{2}}+\bigl{(}A_{k}+2p_{2}\sin(2z)\bigr{)}\tilde{v}_{\bm{k}}=0,$ (21)
where now $z=Mt$ and
@@
Figure 3: Floquet exponents for (a) T-model and (b) E-model as a function of physical wavenumbers and field amplitudes. Red dashed lines represent the instability scale (24) and continuous cyan lines the corresponding instability scales for the full expansion (4). For the T-model, this last is given by (25), whereas for the E-model is obtained numerically. White dotted lines mark the evolution of the physical $k$-modes. The black region represents the points where $\Re(\mu_{k}^{(T,E)})=0$.
@@
However, this approximation is not accurate for some potentials, even if they behave as quadratic at first order, such as the case of $\alpha$-attractors with $\alpha\ll 1$. Fig. 3 corroborates this fact. It shows the Floquet exponent for both the T- and E-models and compares, in each case, $l_{\text{inst}}^{(Q)}$ (dashed red) with the corresponding spatial scale $l_{\text{inst}}^{(T,E)}$ (cyan) of the full expansion (4). For the T-model, imposing again that $\Re\left(\mu_{k}^{(T)}\right)>0$ in (20) we obtain
$l_{\text{inst}}^{(T)}=\left(\frac{1}{2}\sqrt{-(4M^{2}+6M^{2}\Phi_{1}^{2}+6\lambda\Phi_{1}^{2})\epsilon^{2}+\sqrt{24M^{4}\Phi_{1}^{2}\epsilon^{2}+9(M^{2}-\lambda)^{2}\Phi_{1}^{4}\epsilon^{4}}}\,\right)^{-1}.$ (25)
@@
Looking at the expression for $\mu_{k}^{(E)}$, eqn. (18), one can observe the difficulty in obtaining an analytical expression for the instability scale in this case. For that reason, in Fig. 3(b), $l_{\text{inst}}^{(E)}$ is obtained numerically. White dotted lines represent the evolution of the physical modes, and at the moment those modes cross the instability scale, they start to amplify. These crossing points are different if one uses the full expansion (4) or just the quadratic term, pointing out the importance of considering higher-order terms. Moreover, the enhancement of perturbations can be physically explained with the negativity of the coefficients of the expansion, $\lambda_{3}$ or $\lambda_{T}$. If we treat them as self-interactions of the field, a negative coefficient implies an attractive force, which induces a negative pressure and makes some modes unstable [7, 8]. Nonetheless, as the field decays, the cubic and quartic terms of the expansion (4) became less dominant and one recovers the usual preheating scenario [6, 10, 11, 35] where the potential is well approximated by a parabola and $\tilde{v}_{\bm{k}}\sim a^{3/2}$. This can be seen in Fig. 3(a), where for small amplitudes of the field $l_{\text{inst}}^{(Q)}\sim l_{\text{inst}}^{(T)}$. For Fig. 3(b), one should go to smaller amplitudes to have this parallelism. Lastly, it is important to note that not all perturbations with $\left(\frac{k}{a}\right)^{-1}>l_{\text{inst}}$ undergo amplification. The upper limit is set by the Hubble radius $R_{H}=H^{-1}$ and the set of modes defined by $R_{H}>\left(\frac{k}{a}\right)^{-1}>l_{\text{inst}}$ is usually called instability band (IB). Figs. 4(a) and 4(b) show the IB (blue shaded area) for a T- and an E-model, respectively. We see again that, for the T-model, $l_{\text{inst}}^{(T)}$ approaches $l_{\text{inst}}^{(Q)}$ sooner that in the E-model case. This is mainly because the former one is closer to the parabola, given the absence of the asymmetric coefficient $\lambda_{3}$. If we neglect the higher-order terms in the expansion of the potential then we are not considering the amplification of some modes that enter the IB around the beginning of preheating, indicated by the lower gray dashed lines.
6 Applications
@@
  • Primordial Black Holes: Using general relativity one can study the collapse into a black hole (BH) of a real, minimally coupled, massive scalar field in an asymptotically Einstein-de Sitter spacetime background. In [36, 10, 37] it is shown that density perturbations $\delta\rho_{\bm{k}}$ can collapse into a BH provided that the time the $k$-mode spends inside the IB exceeds the time it needs to collapse, defined as
@@
    where $t_{\text{bc}}(k)$ is the time at which each mode enters the instability band and $\delta_{\bm{k}}=\frac{\delta\rho_{\bm{k}}}{\rho}$ is the density contrast, being $\rho$ the background energy density and $\delta\rho_{\bm{k}}$ the density perturbations. In our specific scenario, perturbations amplify faster with decreasing $\alpha$ values. This suggests that PBHs form more rapidly and consequently with smaller masses. However, a critical threshold should exist where PBHs may become overproduced, thereby establishing a lower bound for $\alpha$. This is something we intend to study and publish soon somewhere [26]
  • Scalar-Induced Gravitational Waves: At second order in perturbation theory, the scalar perturbations couple to the tensor ones, inducing the production of GWs [16, 17, 18, 19, 20]. The resulting differential equation for the tensor perturbations $h_{\bm{k}}(\eta)$ in cosmic time $\eta$ is given by
@@
    where $\mathcal{H}=aH$ is the conformal Hubble rate and $\mathcal{S}(\mathbf{k},\eta)$ is the so-called source function, which directly depends on the metric scalar perturbation $\Phi_{\bm{k}}$. This last can be related to the curvature perturbation $\mathcal{R}_{\bm{k}}$ [35] and thus with $v_{\bm{k}}$, which again suggests that for decreasing values of $\alpha$ the fractional energy density of GWs, $\Omega_{\text{GW}}(k)$, can be high enough to reach the range of detection of some actual and future GWs detectors [38, 39]. The fractional energy density is defined as [18]
@@
    where $\rho_{\text{GW}}$ is the energy density of GWs and $\mathcal{P}_{h,\theta}(k)$ the power spectrum of tensor perturbations, with $\theta$ accounting for both polarization states. We will also study this and publish our findings [27]
  • Oscillons: Oscillons are defined as relatively long-lived and spatially localized regions where a scalar field oscillates with large amplitude [21, 22, 23, 24, 25]. They form when the inflationary potential deviates from parabolic shape and could leave some imprint in the present-day universe, particularly in the form of GWs.
@@ 7 Conclusions
This work introduces an analytical model to explain the amplification of curvature perturbations at small scales during the post-inflationary preheating phase, based on the phenomenon of self-resonance. Initially, we solve perturbatively for the inflaton field and expand the inflationary potential up to fourth order. Then, we transform the MS (11) equation into a Hill’s one (12) and provide an explicit formula (18) for Floquet’s exponents of the MS variable. This, so far in the literature, was not given when the potential includes both cubic and quartic self-interactions. Traditionally, self-resonance has been studied primarily within the framework of the perturbed Klein-Gordon equation. However, our approach is applied to the MS equation, since it is more complete in the sense of describing perturbations. While applicable to any potential that can be expressed as (4), we focus our analysis on $\alpha$-attractor models, given their observational favorability [5]. Particularly, for $\alpha\ll 1$, our model reveals a pronounced self-resonance effect, leading to a rapid and substantial enhancement of the curvature perturbations. This analytical model is compared with a full numerical computation, as can be seen in Fig. 1, showing a good agreement, especially concerning the T-model. Also, we stress the importance of avoiding the parabola approximation during preheating for some models, specifically during the transition from inflation to preheating, as illustrated in Figs. 3 and 4. Moreover, this investigation opens a window on the potential role of $\alpha$-attractors in the formation of PBHs on small scales as well as in the production of SIGWs and oscillons during preheating. Furthermore, it suggests a possible physical lower bound on the parameter $\alpha$, since as we decrease its value, perturbations amplify more drastically, and thus for a very extremely low value of $\alpha$ this could potentially leave an imprint in CMB observations. Finally, we stress that this mechanism of amplification of perturbations is inherent to preheating in $\alpha$-attractor models and does not require any fine-tuning.
[…]
Acknowledgements

### v2

@@ 3 Characterization of self-resonance
Here we present the specific Floquet’s exponents for both T- and E-models and compare them with numerical results. These findings can be generalized to any potential whose expansion around $\phi=0$ matches either the expansion of a T- or an E-model.
Figure 2: Curvature perturbations for (a) E-Model $\alpha=10^{-1.5}$, (b) T-Model $\alpha=10^{-2}$, (c) E-Model $\alpha=10^{-2.5}$ and (d) T-Model $\alpha=10^{-3}$ evaluated 5 e-folds after the end of inflation. Numerical computation is shown in continuous black and Floquet theory as described in this work in dashed green. The vertical grey line marks the scale that exits the horizon at the end of inflation, $k_{\text{end}}$. For comparison, black dotted curves show the numerical computation of curvature perturbations at the end of inflation.
@@ 3.1 T-model
In this case we have that $\lambda_{3}=0$ and thus $q_{1}=p_{1}=0$, which greatly simplifies the expression for the Floquet’s exponent
@@
To check the validity of this approximation, we have computed the curvature perturbations numerically, using (28) with Bunch-Davies initial conditions and (18). Then, we have used the Floquet theory derived here, where the scale factor, needed to parametrize the decay of the field, is obtained by self-consistently solving the Friedmann equation (5b), together with the equation of motion of the field (5a). This is important, since considering that the time dependence of the scale factor follows a matter-dominated power-law, that is, $a\sim t^{2/3}$, is incorrect during the first e-folds of preheating, which is precisely where this works mainly focus. Both methods are evaluated 5 e-folds after the end of inflation. Results are displayed in Figs. 2(b) and 2(d) for $\alpha=10^{-2}$ and $\alpha=10^{-3}$ and it shows a good agreement between both methods. Going to smaller values of $\alpha$ implies entering into the non-linear regime, where perturbations backreact and spoil the linear evolution. See B for further details and a comparison with previous estimations. The peaks to the right of the highest peak in Fig. 2 correspond to the amplification in higher resonance bands. One could study these bands by considering the frequencies $n=\pm 2,\pm 3\dots$ and their corresponding couplings. However, to do so, one has to increase each time the order in $\epsilon$, increasing the number of harmonics that come into play. We find it sufficient for our purposes to focus on the first resonance band, as it is where the amplification is most pronounced. Finally, in Fig. 3(a) we display $\Re(\mu_{k}^{(T)})$ for $\alpha=10^{-2.5}$ as a function of $q_{2}$ and $A_{k}$, as well as the trajectories of some modes in the $(q_{2},A_{k})$ plane. We remark that in Fig. 3(a), $p_{2}=0$ and thus the $\mu_{k}^{(T)}$ shown does not correspond to the real time-evolution. However, it is enough to understand why the peak is produced for scales around $k\simeq 10\,k_{\text{end}}$. Those scales cross the $(q_{2},A_{k})$ plane throughout the maximum of $\mu_{k}^{(T)}$. Also, it is possible to give an analytical estimation of the position of the peak from the Floquet’s exponent (19). For $k<k_{\text{end}}$, all scales show the same evolution in the $(q_{2},A_{k})$ plane. This is the reason why all of them are amplified in the same amount.
Figure 3: (a) Floquet chart for a T-model with $p_{2}=0$ (just for graphical purposes) and $\alpha=10^{-2.5}$. White dashed lines correspond to the evolution of the modes in the $(q_{2},A_{k})$ plane and are, from bottom to top: 1, 10, 15, 20, 30, in units of $k_{\text{end}}$. (b) Evolution of $\Re(\mu_{k,E})$ for an E-model with $\alpha=10^{-2.5}$ as a function of the amplitude of the field and for different values of $k$.
@@ 3.2 E-model
Now we have that $\lambda_{3}\neq 0$ and thus the expression for the Floquet exponent is the full one given in (17), we will call it $\mu_{k}^{(E)}$. Figs. 2(a) and 2(c) display the curvature perturbations computed using the same method as in Figs. 2(b) and 2(d) (see previous section), this time for the E-model and for $\alpha=10^{-1.5}$ and $\alpha=10^{-2.5}$. Despite a numerical factor of $\mathcal{O}(1)$ of discrepancy between the numerics and Floquet theory, we find it enough to explain at which scales occur the maximum amplification, which could be also obtained analytically from the full expression (17). The discrepancy between numerics and the analytical result could be because we are neglecting some higher harmonics, which in the case of the E-model could contribute more significantly than in the T-model, as well as more couplings between them. We have observed analytically and also numerically that the amplification of $\mathcal{R}_{\bm{k}}$ is higher for the E-model compared to a T-model with the same value of $\alpha$. This is due to the asymmetry of the potential since the perturbative solution for the E-model has even harmonics and, particularly, a drift term that appears already at order $\beta^{2}$. This fact makes the Floquet’s exponent (17) have more contributions that enhance the amplification of the perturbations. Additionally, we have also observed that the position of the peak of maximum amplification is at higher $k$ for the E-model, compared to a T-model with the same $\alpha$. Now, regarding the time evolution of $\mu_{k}^{(E)}$ for each $k$-mode, a plot like Fig. 3(a) would not be sufficiently illustrative, given that $\mu_{k}^{(E)}$ depends on 5 parameters. This makes it challenging to understand why the peak is around $k\sim 11\,k_{\text{end}}$. Instead, Fig. 3(b) shows $\Re(\mu_{k}^{(E)})$ for $\alpha=10^{-2.5}$ as a function of the amplitude of the field for different values of $k$, expressed in units of $k_{\text{end}}$. Here, we observe that for scales around $k\simeq 11\,k_{\text{end}}$, the Floquet exponent exceeds that of other $k$-modes, resulting in higher amplification.
@@ 4 Contrast with quadratic approximation
During preheating, the potential is usually approximated by a parabola ($\lambda_{3}=\lambda=0$), where the solution for the field is just given by $\phi\simeq\phi_{\text{end}}\left(\frac{a_{\text{end}}}{a}\right)^{3/2}\cos(\tau)$ Kofman:1997yn ; Jedamzik:2010dq ; Martin:2020fgl ; Martin:2019nuw ; Ballesteros:2024hhq . Considering this and keeping just the dominant terms, (28) is easily reformulated into the following Mathieu equation
$\tilde{v}_{\bm{k}}^{\prime\prime}+\bigl{(}A_{k}+2p_{2}\sin(2z)\bigr{)}\tilde{v}_{\bm{k}}=0,$ (20)
where now $\tau=Mt$ and
@@
Figure 4: Floquet exponents for (a) T-model and (b) E-model with $\alpha=10^{-2.5}$ as a function of physical wavenumbers and field amplitudes. Red dashed lines represent the instability scale (23) and continuous cyan lines the corresponding instability scales for the full expansion (4). For the T-model, this last is given by (24), whereas for the E-model is obtained numerically. White dotted lines mark the evolution of the physical $k$-modes. The black region represents the points where $\Re(\mu_{k}^{(T,E)})=0$.
@@
This is the standard preheating scenario. However, this approximation is not accurate for some potentials, even if they behave as quadratic at first order, such as the case of $\alpha$-attractors with $\alpha\ll 1$. Fig. 4 corroborates this fact. It shows the Floquet exponent for both the T- and E-models and compares, in each case, $l_{\text{inst}}^{(Q)}$ (dashed red) with the corresponding spatial scale $l_{\text{inst}}^{(T,E)}$ (cyan) of the full expansion (4). For the T-model, imposing again that $\Re\left(\mu_{k}^{(T)}\right)>0$ in (19) we obtain
$l_{\text{inst}}^{(T)}=2\Biggl{(}-(4M^{2}+6M^{2}\Phi_{1}^{2}+6\lambda\Phi_{1}^{2})\epsilon^{2}+\sqrt{24M^{4}\Phi_{1}^{2}\epsilon^{2}+9(M^{2}-\lambda)^{2}\Phi_{1}^{4}\epsilon^{4}}\,\Biggr{)}^{-1/2}.$ (24)
Looking at the expression for $\mu_{k}^{(E)}$, eqn. (17), one can observe the difficulty in obtaining an analytical expression for the instability scale in this case. For that reason, in Fig. 4(b), $l_{\text{inst}}^{(E)}$ is obtained numerically. White dotted lines represent the evolution of the physical modes, and at the moment those modes cross the instability scale, they start to amplify. These crossing points are different if one uses the full expansion (4) or just the quadratic term, pointing out the importance of considering higher-order terms. Moreover, the enhancement of perturbations can be physically explained with the negativity of the coefficients of the expansion, $\lambda_{3}$ or $\lambda_{T}$. If we treat them as self-interactions of the field, a negative coefficient implies an attractive force, which induces a negative pressure and makes some modes unstable. Nonetheless, as the field decays, the cubic and quartic terms of the expansion (4) became less dominant and one recovers the usual preheating scenario Kofman:1997yn ; Martin:2019nuw ; Martin:2020fgl ; del-Corral:2023apl where the potential is well approximated by a parabola and $\tilde{v}_{\bm{k}}\sim a^{3/2}$. This can be seen in Fig. 4(a), where for small amplitudes of the field $l_{\text{inst}}^{(Q)}\sim l_{\text{inst}}^{(T)}$. For Fig. 4(b), one should go to smaller amplitudes to have this parallelism. Both of these figures are made using $\alpha=10^{-2.5}$. Lastly, it is important to note that not all perturbations with $\left(\frac{k}{a}\right)^{-1}>l_{\text{inst}}$ undergo amplification. The upper limit is set by the Hubble radius $R_{H}=H^{-1}$ and the set of modes defined by $R_{H}>\left(\frac{k}{a}\right)^{-1}>l_{\text{inst}}$ is usually called instability band (IB). Figs. 5(a) and 5(b) show the IB (blue shaded area) for a T- and an E-model with $\alpha=10^{-2.5}$, respectively. We see again that, for the T-model, $l_{\text{inst}}^{(T)}$ approaches $l_{\text{inst}}^{(Q)}$ sooner that in the E-model case. This is mainly because the former one is closer to the parabola, given the absence of the asymmetric coefficient $\lambda_{3}$. If we neglect the higher-order terms in the expansion of the potential then we are not considering the amplification of some modes that enter the IB around the beginning of preheating, indicated by the lower gray dashed lines, which could lead to wrong estimations.
5 Applications
@@
  • Primordial Black Holes: Using general relativity one can study the collapse into a black hole (BH) of a real, minimally coupled, massive scalar field in an asymptotically Einstein-de Sitter spacetime background. In Goncalves:2000nz ; Martin:2019nuw ; Barroso:2024cgg it is shown that density perturbations $\delta\rho_{\bm{k}}$ can collapse into a BH provided that the time the $k$-mode spends inside the IB exceeds the time it needs to collapse, defined as
@@
    where $t_{\text{bc}}(k)$ is the time at which each mode enters the instability band and $\delta_{\bm{k}}=\frac{\delta\rho_{\bm{k}}}{\rho}$ is the density contrast, being $\rho$ the background energy density and $\delta\rho_{\bm{k}}$ the density perturbations. In our specific scenario, perturbations amplify faster with decreasing $\alpha$ values. This suggests that PBHs form more rapidly and consequently with smaller masses. However, a critical threshold should exist where PBHs may become overproduced, thereby establishing a lower bound for $\alpha$. This is something we intend to study and publish soon somewhere draft-alpha-PBH .
  • Scalar-Induced Gravitational Waves: At second order in perturbation theory, the scalar perturbations couple to the tensor ones, inducing the production of GWs Maggiore:2007 ; Guzzetti:2016 ; Domenech:2021 ; Baumann:2007zm ; Ananda:2006af . The resulting differential equation for the tensor perturbations $h_{\bm{k}}(\eta)$ in cosmic time $\eta$ is given by
@@
    where $\mathcal{H}=aH$ is the conformal Hubble rate and $\mathcal{S}(\mathbf{k},\eta)$ is the so-called source function, which directly depends on the metric scalar perturbation $\Phi_{\bm{k}}$. This last can be related to the curvature perturbation $\mathcal{R}_{\bm{k}}$ del-Corral:2023apl and thus with $v_{\bm{k}}$, which again suggests that for decreasing values of $\alpha$ the fractional energy density of GWs, $\Omega_{\text{GW}}(k)$, can be high enough to reach the range of detection of some actual and future GWs detectors Gehrman:2022imk ; Aggarwal:2020olq . The fractional energy density is defined as Domenech:2021
@@
    where $\rho_{\text{GW}}$ is the energy density of GWs and $\mathcal{P}_{h,\theta}(k)$ the power spectrum of tensor perturbations, with $\theta$ accounting for both polarization states. We will also study this and publish our findings draft-alpha-GWs .
@@ 6 Conclusions
This work introduces an analytical model to explain the amplification of curvature perturbations at small scales during the post-inflationary preheating phase, based on the phenomenon of self-resonance. We consider both symmetric and asymmetric inflationary potentials, i.e., T and E $\alpha$-attractors modelsKallosh:2013daa ; Kallosh:2013hoa ; Kaiser:2013sna ; Kallosh:2015lwa . Initially, we solve perturbatively for the inflaton field and expand the inflationary potential up to fourth order. Then, we transform the MS equation into a Hill’s one (13) (see also A) and provide an explicit formula (17) for Floquet’s exponents of the MS variable. This, so far in the literature, was not given when the potential includes both cubic and quartic self-interactions. In addition, the analytical results allow us to estimate, avoiding time-consuming numerical codes, the magnitude of the amplification of curvature perturbations during preheating, as well as the position of the peak in terms of comoving wavenumber $k$ and the instability scale above which perturbations grow. Traditionally, self-resonance has been studied primarily within the framework of the perturbed equation of motion of the field Kofman:1997yn ; Iarygina:2018kee ; Iarygina:2020dwe ; Krajewski:2018moi ; Lozanov:2017hjm ; Jedamzik:2010dq ; Martin:2019nuw ; Sfakianakis:2018lzf ; Ballesteros:2024hhq ; Martin:2020fgl ; Shafi:2024jig ; Mahbub:2023faw ; Sang:2020kpd ; Zhang:2023hjk ; Child:2013ria . However, our approach is applied to the MS equation, which captures the dynamics of the curvature perturbations. While applicable to any potential that can be expressed as (4), we focus our analysis on $\alpha$-attractor models, given their observational favorability Planck:2018jri . Particularly, for $\alpha\ll 1$, our model reveals a pronounced self-resonance effect, leading to a rapid and substantial enhancement of the curvature perturbations, specifically for asymmetric potentials such as E-models due to the drift term. This analytical model is compared with a full numerical computation, as can be seen in Fig. 2, showing a good agreement, especially concerning the T-model. Also, we stress the importance of avoiding the standard parabola approximation during preheating for some models, specifically during the transition from inflation to preheating, as illustrated in Figs. 4 and 5. Moreover, this investigation opens a window on the potential role of $\alpha$-attractors in the formation of PBHs on small scales as well as in the production of SIGWs and oscillons during preheating. Furthermore, it suggests a possible physical lower bound on the parameter $\alpha$, since as we decrease its value, perturbations amplify more drastically, potentially reaching the non-linear regime (see B) and thus for a very extremely low value of $\alpha$ this could potentially impact on the production of primordial black holes or even scalar-induced gravitational-waves. Finally, we stress that this mechanism of amplification of perturbations is inherent to preheating in $\alpha$-attractor models and does not require any fine-tuning.
Acknowledgments